# Mini-Splatting: Representing Scenes with a Constrained Number of Gaussians

### v1

@@ Abstract
In this study, we explore the challenge of efficiently representing scenes with a constrained number of Gaussians. Our analysis shifts from traditional graphics and 2D computer vision to the perspective of point clouds, highlighting the inefficient spatial distribution of Gaussian representation as a key limitation in model performance. To address this, we introduce strategies for densification including blur split and depth reinitialization, and simplification through Gaussian binarization and sampling. These techniques reorganize the spatial positions of the Gaussians, resulting in significant improvements across various datasets and benchmarks in terms of rendering quality, resource consumption, and storage compression. Our proposed Mini-Splatting method integrates seamlessly with the original rasterization pipeline, providing a strong baseline for future research in Gaussian-Splatting-based works.
@@ 1 Introduction
The optimization of minimal Gaussian representation remains a relatively underexplored topic, with only a handful of compression-focused studies venturing into this area. Drawing inspiration from voxel pruning techniques used in grid-based neural representations [38, 20], Niedermayr et al. [26] proposed pruning Gaussians based on parameter sensitivity derived from image gradients. Similarly, Lee et al. [19] and LightGaussian [10] have implemented pruning strategies based on Gaussian opacity and scale. However, these methods, which primarily aim at storage compression, often neglect the inefficient spatial distribution of input Gaussians. As a result, direct pruning, as practiced in these approaches, tends to yield suboptimal simplification outcomes. Moreover, most existing 3D data simplification algorithms focus on preserving feature points [15, 18] or geometric structures [27, 22], making them less suitable for neural representations like 3DGS. Therefore, adapting these techniques to the 3DGS model presents a non-trivial challenge.
@@
In this paper, we start with a comprehensive analysis of Gaussian representation. By visualizing the projected Gaussian centers and rendered results, we find that 3D Gaussians are not uniformly distributed around the object surface, which limits the rendering speed and quality of the model. To construct an efficient Gaussian representation, we pivot the task of directly simplifying 3D Gaussians to reorganizing their spatial positions employing our Gaussian densification and simplification method. Specifically, our proposed densification algorithm includes two key components: blur split, which involves splitting Gaussians in the blurring areas, and depth reinitialization, which uses merged depth points to reinitialize the scene. Our simplification approach comprises binarization, which preserve Gaussians with maximal impact on the rendered image, and sampling, designed to preserve both the geometric structure and rendering quality of the original model. By stacking our densification and simplification techniques, we achieve a more uniform and efficient spatial distribution of Gaussians, as illustrated in Fig. 1. Our main contributions are as follows:
@@
  • Through an in-depth analysis of the spatial distribution of Gaussians, we observe that the vanilla 3DGS presents two significant phenomena: ‘overlapping’ and ‘under-reconstruction’. We identify that this inefficient distribution limits the rendering quality and speed, and it is non-trivial to achieve a minimal Gaussian representation while maintaining rendering quality.
@@
  • We propose a Gaussian densification and simplification algorithm to reorganize the spatial positions of Gaussians rather than directly pruning them. The proposed densification (blur split and depth reinitialization) and simplification (Gaussian binarization and sampling) strategies, leveraging both screen-space and world-space information, effectively achieve a dense distribution of Gaussians around the object.
@@ 2.1 Gaussian Splatting
With recent advancements in neural radiance fields, several works have been proposed to simplify neural representations, primarily focusing on feature grids and introducing voxel pruning [6, 20] or voxel masking [35, 30] strategies. Inspired by these developments, similar pruning techniques have been adapted in recent 3DGS compression-oriented works [26, 19, 10]. For instance, Niedermayr et al. [26] proposed to prune Gaussians based on image gradients, while Lee et al. [19] and LightGaussian [10] performed pruning according to Gaussian opacity and scale. However, these works overlook the issue of inefficient spatial distribution of Gaussians, which can easily lead to suboptimal simplification results after Gaussian pruning.
@@ 3.1 Preliminaries of Gaussian Splatting
Gaussian Representation. 3D Gaussian Splatting (3DGS) [16] models the scene geometry as a set of 3D Gaussians $\{\mathcal{G}_{i}|i=1,\cdots,N\}$. Each Gaussian $\mathcal{G}_{i}$ is characterized by its opacity $\alpha_{i}\in[0,1]$, center $\mathbf{p}_{i}\in\mathbb{R}^{3\times 1}$ and covariance matrix in world space $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ [40] as $\mathcal{G}_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{i})^{T}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\mathbf{p}_{i})}$. To render an image from these Gaussians, 3DGS sorts them in an approximate depth order based on the distance from their centers $\mathbf{p}_{i}$ to the image plane, and further applies applies alpha-blending [23] as follows:
@@ 4.1 Densification
Here, we present our formulation of the rendered depth. Given one elliptical Gaussian defined by its scale $\mathbf{s}=(s_{x},s_{y},s_{z})$, we model the Gaussian as an ellipsoid in 3D space as $g(x,y,z)=\frac{x^{2}}{s_{x}^{2}}+\frac{y^{2}}{s_{y}^{2}}+\frac{z^{2}}{s_{z}^{2}}=1$. With an input ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ and $\mathbf{d}$ are the ray origin and the ray direction transferred in the coordinate system of the ellipsoid, we can simply calculate the position of the mid-point between two intersection points of the ray and the ellipsoid. We set the depth of this mid-point as our Gaussian depth $d^{mid}_{i}$. Another intuitive formulation is to calculate the optimized $t$ where the maximum density of the Gaussian distribution along the ray is reached. We adopt the former formulation because these two share the same numerical result, but the former one can provide an auxiliary discriminant (i.e., the discriminant to distinguish the intersection point of the ray and the ellipsoid exists or not). Details of the calculation are provided in the appendix. For multiple Gaussians, we only collect the Gaussian with the maximum contribution to the pixel to avoid the artifacts from alpha blending. The final depth $d^{mid}$ can be formulated as
$d^{mid}=d^{mid}_{i_{max}},\quad\text{where }i_{max}=\arg\max\limits_{i}w_{i}.$ (3)
In practice, we observe that setting the depth based on the Gaussian center, as $d^{center}=d_{i_{max}}$, can also achieve a similar rendering quality as $d^{mid}$ after further optimization. However, considering that $d^{mid}$ shows a better reconstruction result as illustrated in Figs. 5 (a) and (b), we chose to include it in our main pipeline of Mini-Splatting for potential future usage.
@@ 4.2 Simplification
After the densification process, we achieve a denser Gaussian representation of the scene. To further refine this representation, we propose simplification techniques, including binarization and sampling, to attain an efficient Gaussian representation while still maintaining high-quality rendering.
Gaussian Binarization. Drawing inspiration from the concept of ray-mesh intersection, our Gaussian binarization technique is designed to discard Gaussians that do not directly intersect with the ray. However, we avoid strictly converting the smooth opacity or blending weights into binary values of ${0,1}$, as this could potentially compromise the rendering quality.
In the depth reinitialization step, we model 3D Gaussians as ellipsoids and determine their depths based on the mid-points. We render the depth of the Gaussian with the maximum contribution, representing it as the depth of the scene. These mid-points, indicating the intersection of the rays with the scene representation, serve as the basis for our binarized Gaussians $\mathcal{G}^{bin}$. Specifically, for each image, we only retain Gaussians that provide these intersection points, and formulate these Gaussians as the binarized Gaussians:
$\mathcal{G}^{bin}=\{\mathcal{G}_{i}|i\in\mathcal{I}_{max}\land i\in[1,N]\},$ (4)
@@
The concept of pruning in neural representation is initially introduced for grid-based neural radiance fields [38, 20], aimed at removing empty or low-information voxels, and this technique can also be extended to Gaussian representations. By assigning each Gaussian a pre-defined importance value $I_{i}$, we can simply prune Gaussians with low importance. While this strategy is effective for removing a limited number of Gaussians, it may lead to a collapse in rendering quality when the pruning ratio is relatively large. The main issue lies in the fact that Gaussian importance only reflects its significance before simplification. Direct pruning can easily disrupt local geometry, which in turn degrades the final rendering quality after further optimization. Specifically, neighboring Gaussians often exhibit similar importance in a given area, causing them to be either removed or preserved simultaneously, thus risking the destruction of local geometry. To illustrate this point, we conduct Gaussian pruning on our binarized Gaussians and measure the geometry quality using the chamfer distance between the centers of pruned and non-simplified Gaussians. The significant chamfer distance observed after pruning, especially at low preserving ratios, suggests that relying solely on Gaussian importance leads to suboptimal simplification results.
@@ 6 Experiments
Mini-Splatting-D 0.832 27.54 0.175 4.69 0.854 23.29 0.142 4.26 0.907 29.99 0.210 4.62
Mini-Splatting 0.822 27.30 0.217 0.49 0.836 23.29 0.203 0.20 0.908 29.97 0.253 0.35
@@ 6.1 Experimental Results
The rendered images are are presented in Fig. 8. These qualitative outcomes are in line with the quantitative results provided in Table 1 and Fig. 7. Notably, our Mini-Splatting-D exhibits superior rendering quality in terms of finer details, while our Mini-Splatting maintains an acceptable rendering quality despite employing a limited number of Gaussians. These results indicate the effectiveness of our densification and simplification algorithm.
@@ 6.2 Ablation Studies
Simplification. The ablations of simplification on Mip-NeRF360 are depicted in Fig. 10. Starting from the 3DGS baseline, we first add densification to obtain our Mini-Splatting-D. Subsequently, we integrate direct pruning into Mini-Splatting-D and tune the pruning ratio to establish the baseline curve, Add Pruning, for our simplification algorithm. It is worth noting that this baseline encompasses the combination of Mini-Splatting-D and direct pruning, surpassing 3DGS with pruning (i.e., VQ) as depicted in Fig. 7. To underscore the effectiveness of our Gaussian binarization and sampling techniques, we add binarization and then sampling to Mini-Splatting-D (Add Binarization and Add Sampling). The results clearly demonstrate that our simplification algorithm outperforms direct pruning, achieving approximately a halved reduction in the number of Gaussians while maintaining comparable rendering quality. This result indicates the effectiveness of the proposed simplification method.
@@ 7 Conclusion
We introduced Mini-Splatting, which aimed at representing scene with a constrained nubmer of Gaussians, through our proposed densification and simplification algorithm. Our Gaussian densification process, involving blur splitting and depth initialization, facilitates the construction of a dense Gaussian representation, while our simplification algorithm, featuring Gaussian binarization and sampling, effectively limits the number of Gaussians while preserving high-quality rendering. Our experimental results show the effectiveness of Mini-Splatting and its variants across benchmarks, encompassing not only rendering quality but also resource consumption and storage compression.
@@ Appendix 0.F Details of Mini-Splatting and its Variants
15:              Binarization() $\triangleright$ Our Gaussian Binarization
@@
20:              Binarization() $\triangleright$ Our Gaussian Binarization
@@ Appendix 0.H Additional Quantitative Results
Mini-Splatting-D 0.797 0.642 0.879 0.805 0.643 0.929 0.913 0.932 0.948 0.889 0.819 0.905 0.909
Mini-Splatting 0.772 0.626 0.848 0.805 0.653 0.922 0.904 0.927 0.938 0.873 0.799 0.904 0.912
@@
Mini-Splatting-D 25.51 21.50 27.68 27.12 22.32 31.67 28.79 31.27 31.99 25.37 21.20 29.42 30.56
Mini-Splatting 25.21 21.53 26.87 27.21 22.68 31.29 28.43 31.33 31.17 25.06 21.51 29.48 30.46
@@
Mini-Splatting-D 0.159 0.255 0.090 0.168 0.261 0.187 0.170 0.116 0.174 0.102 0.182 0.218 0.202
Mini-Splatting 0.225 0.327 0.149 0.199 0.314 0.211 0.199 0.129 0.201 0.160 0.245 0.257 0.250

### v2

@@ Abstract
In this study, we explore the challenge of efficiently representing scenes with a constrained number of Gaussians. Our analysis shifts from traditional graphics and 2D computer vision to the perspective of point clouds, highlighting the inefficient spatial distribution of Gaussian representation as a key limitation in model performance. To address this, we introduce strategies for densification including blur split and depth reinitialization, and simplification through intersection preserving and sampling. These techniques reorganize the spatial positions of the Gaussians, resulting in significant improvements across various datasets and benchmarks in terms of rendering quality, resource consumption, and storage compression. Our Mini-Splatting integrates seamlessly with the original rasterization pipeline, providing a strong baseline for future research in Gaussian-Splatting-based works. Code is available.
@@ 1 Introduction
The optimization of minimal Gaussian representation remains a relatively underexplored topic, with only a handful of compression-focused studies venturing into this area. Drawing inspiration from voxel pruning techniques used in grid-based neural representations [39, 20], Niedermayr et al. [26] proposed pruning Gaussians based on parameter sensitivity derived from image gradients. Similarly, Lee et al. [19] and LightGaussian [10] have implemented pruning strategies based on Gaussian opacity and scale. However, these methods, which primarily aim at storage compression, often neglect the inefficient spatial distribution of input Gaussians. As a result, direct pruning, as practiced in these approaches, tends to yield suboptimal simplification outcomes. Moreover, most existing 3D data simplification algorithms focus on preserving feature points [15, 18] or geometric structures [28, 22], making them less suitable for explicit representations with densely optimizable parameters like 3DGS. Therefore, adapting these techniques to the 3DGS model presents a non-trivial challenge.
@@
In this paper, we start with a comprehensive analysis of Gaussian representation. By visualizing the projected Gaussian centers and rendered results, we find that 3D Gaussians are not uniformly distributed around the object surface, which limits the rendering speed and quality of the model. To construct an efficient Gaussian representation, we pivot the task of directly simplifying 3D Gaussians to reorganizing their spatial positions employing our Gaussian densification and simplification method. Specifically, our proposed densification algorithm includes two key components: blur split, which involves splitting Gaussians in the blurring areas, and depth reinitialization, which uses merged depth points to reinitialize the scene. Our simplification approach comprises intersection preserving, which preserve Gaussians with maximal impact on the rendered image, and sampling, designed to preserve both the geometric structure and rendering quality of the original model. By stacking our densification and simplification techniques, we achieve a more uniform and efficient spatial distribution of Gaussians, as illustrated in Fig. 1. Our main contributions are as follows:
@@
  • Through a deep analysis of the spatial distribution of Gaussians, we observe that the vanilla 3DGS presents two significant phenomena: ‘overlapping’ and ‘under-reconstruction’. We identify that this inefficient distribution limits the rendering quality and speed, and it is non-trivial to achieve a minimal Gaussian representation while maintaining rendering quality.
@@
  • We propose a Gaussian densification and simplification algorithm to reorganize the spatial positions of Gaussians rather than directly pruning them. The proposed densification (blur split and depth reinitialization) and simplification (intersection preserving and Gaussian sampling) strategies, leveraging both screen-space and world-space information, effectively achieve a dense distribution of Gaussians around the object.
@@ 2.1 Gaussian Splatting
With recent advancements in neural radiance fields, several works have been proposed to simplify neural representations, primarily focusing on feature grids and introducing voxel pruning [6, 20] or voxel masking [36, 31] strategies. Inspired by these developments, similar pruning techniques have been adapted in recent 3DGS compression-oriented works [26, 19, 10]. For instance, Niedermayr et al. [26] proposed to prune Gaussians based on image gradients, while Lee et al. [19] and LightGaussian [10] performed pruning according to Gaussian opacity and scale. However, these works overlook the issue of inefficient spatial distribution of Gaussians, which can easily lead to suboptimal simplification results after Gaussian pruning. One concurrent work, RadSplat [27], also share a similar technical pipeline with our algorithm.
@@ 3.1 Preliminaries of Gaussian Splatting
Gaussian Representation. 3D Gaussian Splatting (3DGS) [16] models the scene geometry as a set of 3D Gaussians $\{\mathcal{G}_{i}|i=1,\cdots,N\}$. Each Gaussian $\mathcal{G}_{i}$ is characterized by its opacity $\alpha_{i}\in[0,1]$, center $\mathbf{p}_{i}\in\mathbb{R}^{3\times 1}$ and covariance matrix in world space $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ [41] as $\mathcal{G}_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{p}_{i})^{T}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\mathbf{p}_{i})}$. To render an image from these Gaussians, 3DGS sorts them in an approximate depth order based on the distance from their centers $\mathbf{p}_{i}$ to the image plane, and further applies alpha-blending [23] as follows:
@@ 4.1 Densification
Here, we present our formulation of the rendered depth. Given one elliptical Gaussian defined by its scale $\mathbf{s}=(s_{x},s_{y},s_{z})$, we model the Gaussian as an ellipsoid in 3D space as $g(x,y,z)=\frac{x^{2}}{s_{x}^{2}}+\frac{y^{2}}{s_{y}^{2}}+\frac{z^{2}}{s_{z}^{2}}=1$. With an input ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ and $\mathbf{d}$ are the ray origin and the ray direction transferred in the coordinate system of the ellipsoid, we can simply calculate the position of the mid-point between two intersection points of the ray and the ellipsoid. We set the depth of this mid-point as our Gaussian depth $d^{mid}_{i}$. Another intuitive formulation is to calculate the optimized $t$ where the maximum density of the Gaussian distribution along the ray is reached. We adopt the former formulation because these two share the same numerical result, but the former one can provide an auxiliary discriminant (i.e., the discriminant to distinguish the intersection point of the ray and the ellipsoid exists or not). Details of the calculation are provided in the appendix.
For multiple Gaussians, we only collect the Gaussian with the maximum contribution to the pixel to avoid the artifacts from alpha blending. The final depth $d^{mid}$ can be formulated as $d^{mid}=d^{mid}_{i_{max}}$, where $i_{max}=\arg\max\limits_{i}w_{i}$. In practice, we observe that setting the depth based on the Gaussian center, as $d^{center}=d_{i_{max}}$, can also achieve a similar rendering quality as $d^{mid}$ after further optimization. However, considering that $d^{mid}$ shows a better reconstruction result as illustrated in Figs. 5 (a) and (b), we chose to include it in our main pipeline of Mini-Splatting for potential future usage.
@@ 4.2 Simplification
After the densification process, we achieve a denser Gaussian representation of the scene. To further refine this representation, we propose simplification techniques, including intersection preserving and Gaussian sampling, to attain an efficient Gaussian representation while still maintaining high-quality rendering.
Intersection Preserving. Drawing inspiration from the concept of ray-mesh intersection, our intersection preserving technique is designed to discard Gaussians that do not directly intersect with the ray. However, we avoid strictly converting the smooth opacity or blending weights into binary values of ${0,1}$, as this could potentially compromise the rendering quality.
In the depth reinitialization step, we model 3D Gaussians as ellipsoids and determine their depths based on the mid-points. We render the depth of the Gaussian with the maximum contribution, representing it as the depth of the scene. These mid-points, indicating the intersection of the rays with the scene representation, serve as the basis for our intersected Gaussians $\mathcal{G}^{int}$. Specifically, for each image, we only retain Gaussians that provide these intersection points, and formulate these Gaussians as the intersected Gaussians:
$\mathcal{G}^{int}=\{\mathcal{G}_{i}|i\in\mathcal{I}_{max}\land i\in[1,N]\},$ (3)
@@
The concept of pruning in neural representation is initially introduced for grid-based neural radiance fields [39, 20], aimed at removing empty or low-information voxels, and this technique can also be extended to Gaussian representations. By assigning each Gaussian a pre-defined importance value $I_{i}$, we can simply prune Gaussians with low importance. While this strategy is effective for removing a limited number of Gaussians, it may lead to a collapse in rendering quality when the pruning ratio is relatively large. The main issue lies in the fact that Gaussian importance only reflects its significance before simplification. Direct pruning can easily disrupt local geometry, which in turn degrades the final rendering quality after further optimization. Specifically, neighboring Gaussians often exhibit similar importance in a given area, causing them to be either removed or preserved simultaneously, thus risking the destruction of local geometry. To illustrate this point, we conduct Gaussian pruning on our intersected Gaussians and measure the geometry quality using the chamfer distance between the centers of pruned and non-simplified Gaussians. The significant chamfer distance observed after pruning, especially at low preserving ratios, suggests that relying solely on Gaussian importance leads to suboptimal simplification results.
@@ 6 Experiments
Mini-Splatting-D 0.831 27.51 0.176 4.69 0.853 23.23 0.140 4.28 0.906 29.88 0.211 4.63
Mini-Splatting 0.822 27.34 0.217 0.49 0.835 23.18 0.202 0.20 0.908 29.98 0.253 0.35
@@ 6.1 Experimental Results
The rendered images are presented in Fig. 8. These qualitative outcomes are in line with the quantitative results provided in Table 1 and Fig. 7. Notably, our Mini-Splatting-D exhibits superior rendering quality in terms of finer details, while our Mini-Splatting maintains an acceptable rendering quality despite employing a limited number of Gaussians. These results indicate the effectiveness of our densification and simplification algorithm.
@@ 6.2 Ablation Studies
Simplification. The ablations of simplification on Mip-NeRF360 are depicted in Fig. 10. Starting from the 3DGS baseline, we first add densification to obtain our Mini-Splatting-D. Subsequently, we integrate direct pruning into Mini-Splatting-D and tune the pruning ratio to establish the baseline curve, Add Pruning, for our simplification algorithm. It is worth noting that this baseline encompasses the combination of Mini-Splatting-D and direct pruning, surpassing 3DGS with pruning (i.e., VQ) as depicted in Fig. 7. To underscore the effectiveness of our intersection preserving and sampling techniques, we add intersection preserving and then sampling to Mini-Splatting-D (Add Intersection and Add Sampling). The results clearly demonstrate that our simplification algorithm outperforms direct pruning, achieving approximately a halved reduction in the number of Gaussians while maintaining comparable rendering quality. This result indicates the effectiveness of the proposed simplification method.
@@ 7 Conclusion
We introduced Mini-Splatting, which aimed at representing scene with a constrained nubmer of Gaussians, through our proposed densification and simplification algorithm. Our Gaussian densification process, involving blur splitting and depth initialization, facilitates the construction of a dense Gaussian representation, while our simplification algorithm, featuring intersection preserving and Gaussian sampling, effectively limits the number of Gaussians while preserving high-quality rendering. Our experimental results show the effectiveness of Mini-Splatting and its variants across benchmarks, encompassing not only rendering quality but also resource consumption and storage compression.
@@ Appendix 0.F Details of Mini-Splatting and its Variants
15:              Intersection() $\triangleright$ Our Intersection Preserving
@@
20:              Intersection() $\triangleright$ Our Intersection Preserving
@@ Appendix 0.H Additional Quantitative Results
Mini-Splatting-D 0.798 0.642 0.878 0.804 0.640 0.928 0.913 0.934 0.946 0.890 0.817 0.905 0.908
Mini-Splatting 0.772 0.626 0.847 0.806 0.654 0.921 0.904 0.926 0.940 0.874 0.796 0.904 0.912
@@
Mini-Splatting-D 25.55 21.50 27.67 27.11 22.13 31.41 28.72 31.75 31.72 25.43 21.04 29.32 30.43
Mini-Splatting 25.15 21.58 26.84 27.26 22.72 31.39 28.41 31.17 31.57 25.09 21.28 29.47 30.49
@@
Mini-Splatting-D 0.158 0.255 0.090 0.169 0.262 0.190 0.172 0.114 0.175 0.100 0.181 0.218 0.204
Mini-Splatting 0.225 0.327 0.150 0.199 0.312 0.211 0.199 0.130 0.199 0.159 0.245 0.257 0.250